a piece of lumber which is 55 inches long is cut into two parts such that 2 times the larger part will be 12 more than 5 times the smaller part. How large are each of the pieces of lumber?
Answer by ptaylor(2198)

Let x=the smaller part
Then 55-x=the larger part
Now we are told the following:
2(55-x)=5x+12 simplify
110-2x=5x+12 add 2x to and subtract 12 from each side
110-2x+2x-12=5x+2x+12-12 collect like terms
98=7x
x=14 in--------------------smaller part
55-x=55-14=41 in-----------larger part
CK
2 times larger part=2*41=82 in
5 times smaller part=5*14=70 in
70+12=82
82=82
Hope this helps -----ptaylor